During April, the production department of a process manufacturing system completed a number of units of a product and transferred them to finished goods. Of these transferred units, 84,000 were in process in the production department at the beginning of April and 336,000 were started and completed in April. April's beginning inventory units were 80% complete with respect to materials and 20% complete with respect to conversion. At the end of April, 106,000 additional units were in process in the production department and were 80% complete with respect to materials and 30% complete with respect to conversion. 1.Compute the number of units transferred to finished goods. Units transferred out= ? 2. Compute the number of equivalent units with respect to both materials used and conversion used in the production department for April using the weighted-average method.
Answer #1

Answer

  • Requirement 1
    Units transferred out = 84000 + 336000 = 420,000
  • Requirement 2

Equivalent units for:
>Material = (420000 transferred unit x 100%) + (106000 ending WIP units x 80%) = 420000 + 84800 = 504,800
>Conversion Cost= (420000 transferred unit x 100%) + (106000 ending WIP units x 30%) = 420000 + 31800 = 451,800